OpenStudy (anonymous):

find all unique zeros of the function f(x)=8x2-64x2+160x-128

OpenStudy (anonymous):

did u try it to equalize with 0?

OpenStudy (anonymous):

-56x^2+160x-128=0

OpenStudy (anonymous):

-14x^2+40x-32=0 -7x^2+20x-16=0